An insurance company checks police records on 590 accidents selected at random and notes that teenagers were at the wheel in 90 of them.

a) Construct the 95% confidence interval for the percentage of all auto accidents that involve teenage drivers.

95% Cl= (....%, ....%)
Answer #1

Solution:- Given that sample n = 590, x = 90

p = x/n = 90/590 = 0.1525

q = 1-p = 1-0.1525 = 0.8475

95% confidence for the Z = 1.645

a) 95% confidence interval for the percentage of all auto accidents = p +/- Z*sqrt(p*q/n)

= 0.1525 +/- 1.96*sqrt(0.1525*0.8475/590)

= (0.1235 , 0.1815)

= (12.35% , 18.15%)